There are two groups of birds, one bird ask one group "how many members are you in a group?" they replied "our number is half of the other group and the total number of members in two group plus you is 100.
question:
1. how many birds are there in one group?
2. how many birds are there in onother group?

Answer by vertciel(183)

Hello there,
I recategorised your question.
Let's call:
Bird Group #1 = x
Bird Group #2 = y
If Bird Group #2's number is half of the other group, that means y = 0.5x.
If the total number of members in two groups plus the asking bird is 100, this means x + y + 1 = 100
So we have the following equations which we must solve to find the amount of birds:
y = 0.5x ---------(1)
x + y + 1 = 100 --(2)
Let's simplify (2): x + y = 99
Since we know y = 0.5x, we can substitute y into (2):
x + 0.5x = 99
0.5x = 99
x = 198
Therefore, Group 1 has 198 birds and Group 2 has 0.5x birds which is 99 birds.